OpenStudy (anonymous):

radius=7 circumference=? area=?

OpenStudy (anonymous):

c=2pir a=pir^2

OpenStudy (anonymous):

c=7*2*pi a=7^2(49)pi

OpenStudy (anonymous):

so circumference is 14 pi. area is 49 pi